28. Casey Motors recently reported the following information Net income $600,000. Tax rate = 40%. Interest expense- $200,000. Total invested capital employed = $9 million. -After-tax cost of capital = 10%. What is the companys EVA? a. -198,000 b.-135,000 c. -203,400 d. -180,000 e. -216,000
Answer #1

EVA = NOPAT – (WACC * capital invested)

NOPAT = Operating Income + Interest * (1 - tax rate )

= 600000 + 200000*(1-0.40)

= 600000 + 120000

= 720000

EVA = 720000 - ( Capital Invested * WACC)

= 720000 - (9000000*10%)

= 720000 - 900000 i.e -180000